AthenaSaige:

Walk/Bike Bus Car Row totals Under age 15 60 165 Age 15 and above 65 195 Column totals 152 110 98 360 What percentage of students under age 15 travel to school by car? 11% 18% 41% 80%

AthenaSaige:

axie:

152-65=87(under 15 who walk/ride to school) 87+60=147(number of bus/ride&walk) 165-147= your answer

AthenaSaige:

thank you so much!

axie:

yup have a great day!!

Rylee88:

@axie You cannot give the answer you have to help them. https://questioncove.com/rules

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Rylee88 @axie You cannot give the answer you have to help them. https://questioncove.com/rules \(\color{#0cbb34}{\text{End of Quote}}\) A user can in fact give out an answer as long as they explain why its correct/why its correct. He or She didn't directly state the answer they solved a couple of equations and left out the final one to get the answer. So there was basically no rules broken.